OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x − 3.5 y = −3.5x + 3.5 No solution Infinite solutions One solution Two solutions

pooja195 (pooja195):

Hi there :) We are given 2 equations y = 3.5x − 3.5 <---equation 1 y = −3.5x + 3.5 <----equation 2 Both tell use what y equals to just take equation 1 and plug it into equation 2 3.5x-3.5=-3.5+3.5 Solve for x

OpenStudy (anonymous):

For both equations, solve for x?

OpenStudy (anonymous):

Y =3.5x - 3.5 + 3.5 y + 3.5 = 3.5x _________________ 3.5 3.5 y+3.5 ______ = x 3.5

pooja195 (pooja195):

Solve for only the equation i gave \[\huge~\rm~3.5x-3.5=-3.5x+3.5\]

OpenStudy (anonymous):

3.5x-3.5=-3.5+3.5 3.5+3.5 = 0 3.5x - 3.5 = 0 X Must = 1

OpenStudy (anonymous):

-3.5+3.5 = 0 *

OpenStudy (anonymous):

:D

pooja195 (pooja195):

YES :D Ok now sub in x value to e1 y=3.5x−3.5: y=3.5x−3.5 y=(3.5)(1)−3.5 y=?

OpenStudy (anonymous):

0

pooja195 (pooja195):

So how many solutions are there ?

OpenStudy (anonymous):

... 1?

OpenStudy (anonymous):

or is it 0..

pooja195 (pooja195):

How many pairs of coordinates are there?

OpenStudy (anonymous):

Hmmmmm....................................... ONE

pooja195 (pooja195):

So how many solutions?

OpenStudy (anonymous):

ONE

pooja195 (pooja195):

GOOD :D

OpenStudy (anonymous):

Thanks for the Help.

pooja195 (pooja195):

np

OpenStudy (anonymous):

And hello to you cheaters, looking at this question 5 years in the future!

pooja195 (pooja195):

XD